# Multi-Armed Bandits with Abstention

### 23 Feb 2024 v1

@@ D.1 Upper Bounds
$R^{\mathrm{RW}}_{\mu,c}(T,\pi)$ $=T\cdot\left(\mu_{1}\vee c\right)-\mathbb{E}\left[\sum_{t=1}^{T}\Big{(}X_{t}\cdot\mathbbm{1}\{B_{t}=0\}+c\cdot\mathbbm{1}\{B_{t}=1\}\Big{)}\right]$
$=T\cdot\left(\mu_{1}\vee c\right)-\mathbb{E}\left[\sum_{t=1}^{T}\Big{(}\mu_{A_{t}}\cdot\mathbbm{1}\{B_{t}=0\}+c\cdot\mathbbm{1}\{B_{t}=1\}\Big{)}\right]$
@@
where the penultimate inequality is derived from an argument analogous to that in (6), and the last inequality is a consequence of Lemma B.4.
@@
D.2 Lower Bounds
Proof of Theorem 4.4.
The proof structure for Theorem 4.4 closely parallels that of Theorem 3.5 in Appendix C.2, although certain specific details contain significant variations. Therefore, we will streamline the shared components and elaborate on the distinctions.
@@ D.2 Lower Bounds
However, unlike the asymptotic lower bound part of the proof of Theorem 3.5, we cannot apply the properties of $R^{\mathrm{CA}}$-consistency here, as $R^{\mathrm{RW}}$-consistency does not imply $R^{\mathrm{CA}}$-consistency in general. Instead, we will demonstrate the desired result (21) directly.
@@
Employing a similar approach to the one used in the minimax lower bound part of the proof of Theorem 3.5, utilizing Lemmas B.1 and B.2, we can derive:
@@
The construction employed in the fixed-reward setting here is analogous to the one utilized in the proof of Theorem 3.5.
@@
Similarly, by applying Lemmas B.1 and B.2, we can derive that
@@
Therefore, we have established the instance-independent minimax lower bound.
Appendix E Additional Numerical Results
E.1 Results for the Fixed-Reward Setting
[…]
@@ E.1 Results for the Fixed-Reward Setting
Refer to caption (a) Instance $\mu^{\dagger}$ Refer to caption (b) Instance $\mu^{\ddagger}$
@@
Within each sub-figure, a pattern emerges. As the abstention reward $c$ increases, the empirical average cumulative regret initially remains relatively stable and starts to decline once $c$ crosses a certain threshold, eventually stabilizing around a small value. These observations are consistent with our theoretical expectations. Specifically, when the abstention reward $c$ is lower than the smallest mean reward among the arms, the agent derives no benefit from opting for the abstention action over selecting an arm. On the other hand, when the abstention reward $c$ exceeds the highest mean reward of the arms, abstention becomes the optimal decision and its reward is even superior to choosing the best arm. In this specific scenario, it is possible to achieve a regret of $o(\log T)$; see Remark 4.2 for further insights.
Refer to caption (a) Instance $\mu^{\dagger}$ Refer to caption (b) Instance $\mu^{\ddagger}$
E.2 Random Instances
@@
Refer to caption (a) $K=20$ Refer to caption (b) $K=30$
Refer to caption (a) $K=20$ Refer to caption (b) $K=30$

### 22 Mar 2026 v2

@@ E.1 Upper Bounds
$R^{\mathrm{RW}}_{\mu,c}(T,\pi)$ $=T\cdot\left(\mu_{1}\vee c\right)-\mathbb{E}\left[\sum_{t=1}^{T}\Big(X_{t}\cdot\mathbf{1}\{B_{t}=0\}+c\cdot\mathbf{1}\{B_{t}=1\}\Big)\right]$
$=T\cdot\left(\mu_{1}\vee c\right)-\mathbb{E}\left[\sum_{t=1}^{T}\Big(\mu_{A_{t}}\cdot\mathbf{1}\{B_{t}=0\}+c\cdot\mathbf{1}\{B_{t}=1\}\Big)\right]$
@@
where the penultimate inequality is derived from an argument analogous to that in (6), and the last inequality is a consequence of Lemma 4.
@@
E.2 Lower Bounds
Proof of Theorem 4.
The proof structure for Theorem 4 closely parallels that of Theorem 2 in Appendix D.2, although certain specific details contain significant variations. Therefore, we will streamline the shared components and elaborate on the distinctions.
@@ E.2 Lower Bounds
However, unlike the asymptotic lower bound part of the proof of Theorem 2, we cannot apply the properties of $R^{\mathrm{CA}}$-consistency here, as $R^{\mathrm{RW}}$-consistency does not imply $R^{\mathrm{CA}}$-consistency in general. Instead, we will demonstrate the desired result (22) directly.
@@
Employing a similar approach to the one used in the minimax lower bound part of the proof of Theorem 2, utilizing Lemmas 1 and 2, we can derive:
@@
The construction employed in the fixed-reward setting here is analogous to the one utilized in the proof of Theorem 2.
@@
Similarly, by applying Lemmas 1 and 2, we can derive that
@@
Therefore, we have established the instance-independent minimax lower bound. ∎
Appendix F Numerical Experiments
F.1 Results for the Fixed-Reward Setting
(a) Instance $\mu^{\dagger}$
(b) Instance $\mu^{\ddagger}$
(a) Instance $\mu^{\dagger}$
(b) Instance $\mu^{\ddagger}$
In this subappendix, we present the empirical results pertaining to the fixed-reward setting. Specifically, we examine the empirical performances of two particular realizations of our algorithm FRW-ALGwA (as outlined in Algorithm 2): FRW-TSwA and FRW-UCBwA. The former uses Less-Exploring Thompson Sampling (Jin et al., 2023) as its base algorithm, while the latter employs $\text{KL-UCB}^{++}$ (Ménard & Garivier, 2017). Note that $\text{KL-UCB}^{++}$ is not an anytime algorithm; that is, it requires the prior knowledge of the time horizon $T$ as an input parameter. In the following experiments, we continue to utilize the two bandit instances $\mu^{\dagger}$ and $\mu^{\ddagger}$, as previously defined in Appendix F.1.
In a manner analogous to our methodology for the fixed-regret setting, we adopt the original versions of Less-Exploring TS and $\text{KL-UCB}^{++}$ as baseline algorithms without the abstention option. The experimental results of the different methods with abstention reward $c=0.9$ for different time horizons $T$ are presented in Figure 4. Additionally, we plot the instance-dependent asymptotic lower bound (ignoring the limit in $T$) on the cumulative regret (see Theorem 4) within each sub-figure. From Figure 4, we have the following observations:
@@ F.1 Results for the Fixed-Reward Setting
Within each sub-figure, a pattern emerges. As the abstention reward $c$ increases, the empirical average cumulative regret initially remains relatively stable and starts to decline once $c$ crosses a certain threshold, eventually stabilizing around a small value. These observations are consistent with our theoretical expectations. Specifically, when the abstention reward $c$ is lower than the smallest mean reward among the arms, the agent derives no benefit from opting for the abstention action over selecting an arm. On the other hand, when the abstention reward $c$ exceeds the highest mean reward of the arms, abstention becomes the optimal decision and its reward is even superior to choosing the best arm. In this specific scenario, it is possible to achieve a regret of $o(\log T)$; see Remark 6 for further insights.
F.2 Random Instances
(a) $K=20$
(b) $K=30$
(a) $K=20$
(b) $K=30$